Grade 10Physical Chemistry

In the analysis f 0.500 g sample of feldspar, a mixture of chlorides of sodium and potassium is obtained which weighs 0.1180g. Subsequent treatment of mixed chlorides, with silver nitrate gives 0.2451 g of silver chloride. What is the percentage of sodium oxide and potassium oxide in feldspar.

Hrishant Goswami
12 Years agoGrade 10

1 Answer

Jitender Pal
12 Years ago
(i) Equate given mass of AgCI against mass obtained from NaCI and KCI
(ii) 2NaCI ≡ Na­2O & 2KCI ≡ K2O
Let amount of NaCI in mixture = x gm
∴ amount of KCI in mixture = (0.118 - x) gm
NaCI + AgNO3 → AgCI + NaNO3
58.5 g 143.5 g
∵ 58.5 g NaCI gives AgCI = 143.5 g
∴ x g NaCI gives AgCI = 143.5/58.5 * x g
Again
KCI + AgNO3 → AgCI + KNO3
74.5 g 143.5 g
∵ 74.5 g KCI gives AgCI = 143.5 g
∴ (0.118 - x) g KCI gives AgCI = (143.5/74.5 * 0.118 – x)g
Total weight of AgCI = 0.2451g
∴ (143.5/58.5 * x) + [143.5/74.5 * (0.118 - x)] = 0.2451
∴ x = 0.0338g
∴ Amount of NaCI in mixture = 0.0338g
∴ Amount of KCI in mixture = 0.118 – 0.0338 = 0.0842g
Since 2NaCI ≡ Na2O
2 * 58.5 62
= 117.0
∵ 117g NaCI is equivalent = 62.0g Na2O
∴ 0.033g NaCI is equivalent to = 62.0/117 * 0.0338 g Na2O
= 0.0179g
% of Na2O in 0.5g of feldspar = 0.0179/0.500 * 100 = 3.58%
2KCI ≡ K2O
2 * 774.5 = 149 94
∵ 149g of KCI is equivalent to = 94g K2O
∴ 0.0842g of KCI is equivalent to 94/149 * 0.0842
= 0.0531 g K2O
∴ % of K2O in 0.5 of feldspar = 0.0531/0.5 * 10.62%
% of Na2O in feldspar = 3.58%
% of K2O in feldspar = 10.62%